# Machine Learning in Fundamental Physics Natural Image Classification via Quasi-Cyclic Graph Ensembles and Random-Bond Ising Models at the Nishimori Temperature

### v1

@@ 15 Compute the positive root $\tilde{\beta}=\frac{-b-\sqrt{b^{2}-4ac}}{2a}$;
[…]
@@ 4.7 Graph construction from CNN features
Given an image $x$, the backbone network $f_{\theta}$ (MobileNetV2, [34]) yields a feature vector $\mathbf{h}\in\mathbb{R}^{1280}$. For each class $c$ we retain a subset $\mathcal{I}^{(c)}$ of $s\ll 1280$ most discriminative indices, e.g. those with largest absolute difference between class-conditional means. The reduced feature vector for class $c$ is $\mathbf{z}^{(c)}=(h_{i})_{i\in\mathcal{I}^{(c)}}\in\mathbb{R}^{s}$. A similarity kernel (Cosine, gaussian and etc) builds a weighted adjacency matrix between the $K$ class representatives, (Eq. 8 in [2]):
@@
where $\gamma$ - kernel bandwidth, $d_{\mathrm{cos}}$ - cosine-distance kernel. To enforce sparsity we retain only the $p$ strongest edges per vertex. The resulting sparse adjacency is then mapped onto a MET QC-LDPC parity-check matrix by selecting an appropriate protograph and CPMs (Section 3).
@@ 5 Experimental evaluation
All experiments use the same MobileNetV2 pre-trained backbone, Fig. 5. The feature extractor (CNN layers) is frozen, continuous features are binarized using sign, only the graph embedding (replace of feed-forward MLP) and final classifier are trained. We use following datasets: ImageNet10: 10-class subset (13 000 images, 10 00010\,000 training, 3 0003\,000 testing). ImageNet100: 100-class subset (130 000130\,000 training, 5 0005\,000 testing; 1 3001\,300 training and $50$ test samples per class). The ImageNet10 and ImageNet10 use 32-D spectral and 64-D respectively.
@@
Figure 5: Pipeline: MobileNetV2 feature extraction followed by graph spectral embedding.
5.1 Influence of Topological Invariants on Spectral Embedding
For the purpose of illustration we consider six frequently occurring trapping sets $TS(4,2)$, $TS(4,6)$, $TS(9,2)$, $TS(13,6)$, $TS(26,20)$, $TS(28,22)$. Table 1 collects the most relevant topological and spectral quantities for each of them. From a spectral-embedding perspective a trapping set can be viewed as a locally dense cluster of edges that is attached to only a few low-degree check nodes. Such a configuration produces negative eigenvalues of the Bethe–Hessian matrix $H_{r}$ (see the column “# neg. $\lambda(H_{1})$”) and, more importantly, gives rise to low-energy modes that are poorly separated by any embedding based on the graph Laplacian. Consequently, graphs containing many of these substructures generate highly correlated features, which in turn can dramatically deteriorate the performance of downstream classifiers.
Table 1: Topological and spectral invariants of Trapping sets.
Invariant $TS(4,2)$ $TS(4,6)$ $TS(26,20)$ $TS(9,2)$ $TS(13,6)$ $TS(28,22)$
$\rho$ (spectral radius) 1.618 4.000 2.7545 8.9168 10.4154 13.5644
$r_{\rm crit}=\sqrt{\rho}$ 1.272 2.000 1.6597 2.9861 3.2273 3.6830
$\{\lambda(H_{1})<0\}$ (negative modes at $r=1$) 1 2 1 1 0 0
$K_{0}$ (zero-dimensional Betti number) 1 5 7 12 17 45
$K_{1}$ (one-dimensional Betti number) 1 1 1 0 1 1
Kervaire invariant $\kappa$ 1 1 1 0 0 0
$w_{2}$ (Stiefel–Whitney) 1 1 1 1 1 1
Mod-2 Betti $(\beta_{0},\beta_{1})$ (1,0) (1,0) (1,0) (1,0) (1,0) (1,0)
Bordism obstruction 0 0 0 0 0 1
To assess the practical impact of these invariants we built two graphs that serve as adjacency matrices in a convolutional architecture trained on the ImageNet-100 benchmark. Optimized irregular MET QC-LDPC toric graph. The protograph matrix (3.6 average column weight) is
lifted with a circulant size of $520$, [2]. During the construction we deliberately eliminated harmful trapping sets; the remaining trapping sets is the relatively benign $TS(26,20)$, which possesses a single negative Bethe–Hessian eigenvalue and a modest continuous genus (see Table 1). After a Bayesian classifier we obtained Top-1 accuracy: $0.6724\pm 2.46\times 10^{-6}$; Top-3 accuracy: $0.8582\pm 8.53\times 10^{-7}$. Thus, with an average degree of only $3.6$ we achieve performance comparable to much denser near regular graphs. Non-optimized irregular MET QC-LDPC toric graph (same $\bar{w}=3.6$) was built without filtering out low-genus, high-mode traps. Consequently it contains the sets $TS(4,2)$, $TS(9,2)$, $TS(13,6)$ and $TS(28,22)$. These introduce several additional negative Bethe–Hessian eigenvalues (up to three for $TS(28,22)$) and raise $\widehat{A}$, thereby generating many low-energy directions that corrupt the spectral embedding. The resulting classification scores are Top-1 accuracy: $0.1368\pm 3.72\times 10^{-4}$; Top-3 accuracy: $0.2450\pm 5.63\times 10^{-5}$. The more than four-fold drop in top-1 performance underscores how a few detrimental trapping sets can cripple the discriminative power of graph-based features. For reference we generated dense Erdős–Rényi graphs (no QC structure, original weight 5 and 10) with average degrees $d\approx 3$ and $d\approx 5$ (almost regular) after cycle broken. After six independent runs the observed accuracies were top-1 $0.617467\pm 3.72E-05$ and top-3 $0.8283\pm 2.54E-05$ (for weight 5) and top-1 $0.659333\pm 1.30E-05$ and top-3 $0.85396667\pm 8.4E-06$ (for weight 10). Even though these dense random graphs outperform the non-optimised QC-LDPC design, they are considerably less efficient (higher degree, no structured protection) and lack the flexibility offered by quasi-cyclic constructions. Negative Bethe–Hessian modes identify substructures that generate low-energy directions in the spectral embedding; each such direction reduces feature orthogonality. The continuous genus $\widehat{A}$ provides a quantitative measure of how “dangerous’’ a trap is: larger values correlate with more severe degradation of downstream performance. By pruning high-genus, multi-mode traps (e.g. $TS(4,2)$, $TS(28,22)$) we can keep the graph sparse ($\bar{w}\approx 3.6$) while retaining classification accuracy comparable to much denser Erdős–Rényi graphs, and improve top-3 accuracy important for graph ensemble (using unequal protection of features, introduce high degree columns). The presented design methodology—selecting protographs, eliminating harmful trapping sets, and fine-tuning via Bayesian optimization—yields QC-LDPC graphs that combine structured protection, and strong discriminative power. All auxiliary material (list of small weight TS, TS submatrix, parity-check matrix, eigenvalue trajectories, topology invariants and related quadratic form signatures, source code, etc) is available in the public repository [35].
@@ 5.1 Influence of Topological Invariants on Spectral Embedding
Figure 6: Spectral embedding histogram with adjacency matrix of clusters (visualized by colors) under estimated Nishimori temperature $\beta_{N}$. (left) MET QC-LDPC Torical graph. (right) QC Spherical graph.
5.2 Graph Ensembles for ImageNet-100
@@
We construct two families of MET–QC-LDPC graphs and then combine them into an ensemble that produces a final class decision by majority voting, Fig. 6. Spherical graphs – each instance consists of a single circulant ring, has a fixed column weight $k=10$ and CPM size $L=2600$. Toroidal graphs – each instance is built from two independent circulant rings (forming a torus), uses $L=520$, and exhibits an average column weight of 3.6. Both families are first optimised with respect to the permanent bound (4) and the Bethe-permanent bound [27]. This optimisation explicitly suppresses low-weight trapping sets that would otherwise degrade the spectral embedding. From the two families we generate three affinity graphs (see Fig. 6), feed each graph into an identical classifier and obtain three independent predictions: $\widehat{y}_{1},\;\widehat{y}_{2},\;\widehat{y}_{3}.$ The final label is obtained by a simple majority vote. In the rare case that the three votes do not agree we invoke an arbiter network to resolve the conflict. Formally,
@@
where $\operatorname{Arb}(\cdot)$ denotes the output of the arbiter network. Instead of hard voting we can average the class-wise posterior probabilities supplied by each graph $p_{c}=\frac{1}{3}\sum_{i=1}^{3}p^{(i)}_{c},\widehat{y}_{\text{soft}}=\arg\max_{c}p_{c},$ with $p^{(i)}_{c}$ being the probability of class $c$ estimated by the $i$-th graph. The soft-voting rule often yields smoother confidence scores while preserving the benefits of ensemble diversity.
Table 2: Classification results on ImageNet10 using graph embeddings.
@@ 5.3 Training protocol
For each class we compute $\mathbf{z}^{(c)}$, build a sparse similarity graph, embed it via the Bethe–Hessian at the estimated Nishimori temperature $\beta_{N}$ (Algorithm 1), and extract the $r$ eigenvectors ($r=32$ for ImageNet10, $r=64$ for ImageNet100), Fig. 6 . These eigenvectors constitute a low-dimensional embedding $\mathbf{e}_{i}$ for every training sample. A linear classifier (softmax) is trained on $\{\mathbf{e}_{i}\}$ using cross-entropy loss.
@@
Figure 7: (Left) Confusion matrix heatmap for ImageNet100. (Right) Per-class top-1 accuracy heatmap.
Figure 8: Spectral embedding histogram for ’cock’ and ’hen’ classes
5.4 Results
@@
Table 2 reports top-1 accuracy, precision, recall and $F_{1}$ scores for ImageNet10. Spherical Graph Spectral Embedding achieves $98.7\%$ top-1 accuracy on ImageNet10, while Fig. 7 shows the confusion matrix for ImageNet100. The mixed ensemble combining spherical and toroidal graphs through majority voting yields the best performance. For ten classes, discriminative features rarely overlap, but scaling to one hundred classes causes many informative dimensions to intersect, reducing accuracy for both plain CNNs and spectral classifiers. Our experiments identified optimal graphs for ImageNet100 consisting of a spherical adjacency with weight $k=10$ and circulant size $L=2600$, paired with a toroidal graph having weight $k=10$, protograph size $26\times 26$ (containing 5 non-zero MET CPMs of weight 2), and CPM size $L=100$, as visualized in Fig. 6. This configuration delivers $77.84\%$ top-1 accuracy when processed through the spectral pipeline. To address frequent confusion between the two highest class probabilities, we introduced a lightweight MLP arbitrator trained on the same embeddings. The arbitrator achieves $98.43\%$ average pairwise accuracy on training data, though the most challenging pairs remain at approximately $67\%$ accuracy. During inference, the system examines the three largest spectral scores $\{p_{(1)},p_{(2)},p_{(3)}\}$, invoking the arbitrator when the margin $p_{(1)}-p_{(2)}$ falls below a preset threshold $T$. This uncertainty-driven fallback mechanism boosts top-1 accuracy from $77.84\%$ to $\mathbf{82.73\%}$ while maintaining the original runtime efficiency and memory footprint of the single-graph approach. The Erdős–Rényi graph spectral embedding demonstrates $69.1\%\pm 2.88\times 10^{-5}$ precision under Nishimori temperature across five trials, averaged over all random seeds. Our three-graph ensemble achieves $\mathbf{82.7\%}$ accuracy on ImageNet100 while using just one twentieth of the original feature dimension and memory footprint, representing a $13.6\%$ improvement over standard Erdős–Rényi graphs without large weights. Classification challenges emerge most prominently between class 7 ("cock") and class 8 ("hen"), as shown in Fig. 7, with training accuracy reaching $85\%$ but validation accuracy dropping to $63\%$. The spectral embedding histogram under Nishimori temperature (Fig. 8) reveals Gaussian intersections with heavy tails in this problematic region. Another challenging area occurs around class 44 ("spider"), where we observe three-class overlap in the heatmap visualization, Fig. 7.
@@ 6 Future Directions
The present pipeline keeps the CNN backbone frozen, employs a similarity kernel and uses a MET-QC-LDPC adjacency. A natural next step is to make all three components differentiable and train them jointly. One can back-propagate through the feature extractor $f_{\theta}$, learn a Mahalanobis (or other) metric for the similarity kernel, and adjust the CPM that define the QC-LDPC parity-check matrix by gradient descent or reinforcement learning. Rich multimodal embeddings such as contrastive language-image pre-training provide powerful initializations for this end-to-end scheme. Embedding topological information directly into the loss is another promising avenue. TS correspond to unwanted $\mathbb{Z}/2$-torsion and low-genus cycles, which appear as negative eigenvalues of the Bethe–Hessian. Adding a differentiable penalty proportional to the number (or an estimator) of such negative modes—e.g. could suppress TS during training instead of removing them post-hoc. Extending experiments from ImageNet-100 to the ImageNet-10k benchmark, applying the method to video streams could reveal new ways of handling high-dimensional structured data while preserving a tiny memory footprint suitable for edge devices. The sparse adjacency matrices produced by our design are attractive priors for GNN. Incorporating them into message-passing GCN/GAT layers, using them to sparsify self-attention in Transformers, or employing diffusion-based graph convolutions may combine the interpretability of physics-inspired graphs with the expressive power of learned GNNs.

### v2

@@ 15 Compute the positive root $\tilde{\beta}=\frac{-b-\sqrt{b^{2}-4ac}}{2a}$;
20  Perform one Newton correction: $g=\frac{\lambda_{\min}\!\bigl(B_{\tilde{\beta}+\delta}\bigr)-\lambda_{\min}\!\bigl(B_{\tilde{\beta}}\bigr)}{\delta},\qquad\beta_{i+1}=\tilde{\beta}-\frac{\lambda_{\min}(B_{\tilde{\beta}})}{g}$;
@@ 3.2 Graph construction from CNN features
Given an image $x$, the backbone network $f_{\theta}$ (MobileNetV2, [21]) yields a feature vector $\mathbf{h}\in\mathbb{R}^{1280}$. For each class $c$ we retain a subset $\mathcal{I}^{(c)}$ of $s\ll 1280$ most discriminative indices, e.g. those with largest absolute difference between class‑conditional means. The reduced feature vector for class $c$ is $\mathbf{z}^{(c)}=(h_{i})_{i\in\mathcal{I}^{(c)}}\in\mathbb{R}^{s}$. A similarity kernel (Cosine, gaussian and etc) builds a weighted adjacency matrix between the $K$ class representatives, (Eq. 8 in [7]):
@@
where $\gamma$ - kernel bandwidth, $d_{\mathrm{cos}}$ - cosine‐distance kernel. To enforce sparsity we retain only the $p$ strongest edges per vertex. The resulting sparse adjacency is then mapped onto a MET QC‑LDPC parity‑check matrix by selecting an appropriate protograph and CPMs (Section 3).
@@ 4 Experimental evaluation
All experiments use the same MobileNetV2 backbone pre‑trained on ImageNet, Fig. 4. The feature extractor (CNN layers) is frozen; only the graph embedding (replace of feed-forward MLP) and final classifier are trained. We use following datasets: ImageNet10: 10‑class subset (13 000 images, 10 00010\,000 training, 3 0003\,000 testing). ImageNet100: 100‑class subset (130 000130\,000 training, 5 0005\,000 testing; 1 3001\,300 training and $50$ test samples per class). The ImageNet10 and ImageNet10 use 32-D spectral and 64-D respectively.
@@
Рис. 4: Pipeline: MobileNetV2 feature extraction followed by graph spectral embedding.
@@
Рис. 5: Spectral embedding histogram with adjacency matrix of clusters (visualized by colors) under estimated Nishimori temperature $\beta_{N}$. (left) MET QC-LDPC Torical graph. (right) QC Spherical graph.
4.1 Graph Ensembles for ImageNet‑100
@@
We construct two families of MET–QC‑LDPC graphs and then combine them into an ensemble that produces a final class decision by majority voting, Fig. 5. Spherical graphs – each instance consists of a single circulant ring, has a fixed column weight $k=10$ and CPM size $L=2600$. Toroidal graphs – each instance is built from two independent circulant rings (forming a torus), uses $L=520$, and exhibits an average column weight of 3.6. Both families are first optimized with respect to the permanent bound (4) and the Bethe‑permanent bound [22]. This optimization explicitly suppresses low‑weight trapping sets that would otherwise degrade the spectral embedding. From the two families we generate three affinity graphs (see Fig. 5), feed each graph into an identical classifier and obtain three independent predictions: $\widehat{y}_{1},\;\widehat{y}_{2},\;\widehat{y}_{3}.$ The final label is obtained by a simple majority vote. In the rare case that the three votes do not agree we invoke an arbiter network to resolve the conflict. Formally hard ensemble,
@@
where $\operatorname{Arb}(\cdot)$ denotes the output of the arbiter network. Instead of hard voting we can average the class‑wise posterior probabilities supplied by each graph $p_{c}=\frac{1}{3}\sum_{i=1}^{3}p^{(i)}_{c},\widehat{y}_{\text{soft}}=\arg\max_{c}p_{c},$ with $p^{(i)}_{c}$ being the probability of class $c$ estimated by the ii‑th graph. The soft‑voting rule often yields smoother confidence scores while preserving the benefits of ensemble diversity.
Таблица 1: Classification results on ImageNet10 using graph embeddings.
@@ 4.2 Training protocol
For each class we compute $\mathbf{z}^{(c)}$, build a sparse similarity graph, embed it via the Bethe–Hessian at the estimated Nishimori temperature $\beta_{N}$ (Algorithm 1), and extract the $r$ eigenvectors ($r=32$ for ImageNet10, $r=64$ for ImageNet100), Fig. 5 . These eigenvectors constitute a low‑dimensional embedding $\mathbf{e}_{i}$ for every training sample. A linear classifier (softmax) is trained on $\{\mathbf{e}_{i}\}$ using cross‑entropy loss.
@@
Рис. 6: (Left) Confusion matrix heatmap for ImageNet100. (Right) Per‑class top‑1 accuracy heatmap.
Рис. 7: Spectral embedding histogram for ’cock’ and ’hen’ classes
4.3 Results and Discussion
@@
Table 1 reports top-1 accuracy, precision, recall and $F_{1}$ scores for ImageNet10. Spherical Graph Spectral Embedding achieves $98.7\%$ top-1 accuracy on ImageNet10, while Fig. 6 shows the confusion matrix for ImageNet100. The mixed ensemble combining spherical and toroidal graphs through majority voting yields the best performance. For ten classes, discriminative features rarely overlap, but scaling to one hundred classes causes many informative dimensions to intersect, reducing accuracy for both plain CNNs and spectral classifiers. Our experiments identified optimal graphs for ImageNet100 consisting of a spherical adjacency with weight $k=10$ and circulant size $L=2600$, paired with a toroidal graph having weight $k=10$, protograph size $26\times 26$ (containing 5 non-zero MET CPMs of weight 2), and CPM size $L=100$, as visualized in Fig. 5. This configuration delivers $77.84\%$ top-1 accuracy when processed through the spectral pipeline. To address frequent confusion between the two highest class probabilities, we introduced a lightweight MLP arbitrator trained on the same embeddings. The arbitrator achieves $98.43\%$ average pairwise accuracy on training data, though the most challenging pairs remain at approximately $67\%$ accuracy. During inference, the system examines the three largest spectral scores $\{p_{(1)},p_{(2)},p_{(3)}\}$, invoking the arbitrator when the margin $p_{(1)}-p_{(2)}$ falls below a preset threshold $T$. This uncertainty-driven fallback mechanism boosts top-1 accuracy from $77.84\%$ to $\mathbf{82.73\%}$ while maintaining the original runtime efficiency and memory footprint of the single-graph approach. The Erdős–Rényi graph spectral embedding demonstrates $69.1\%\pm 2.88\times 10^{-5}$ precision under Nishimori temperature across five trials, averaged over all random seeds. Our three-graph hard ensemble achieves $\mathbf{82.7\%}$ accuracy on ImageNet100 while using just one twentieth of the original feature dimension and memory footprint, representing a $13.6\%$ improvement over standard Erdős–Rényi graphs without large weights. Classification difficulties are most evident between class 7 (“cock”) and class 8 (“hen”), as shown in Fig. 6. Although the network reaches a training accuracy of $85\%$ on these two categories, validation performance collapses to $63\%$, indicating severe over‑fitting. The spectral embedding histogram obtained at the Nishimori temperature (Fig. 7) reveals that the corresponding Gaussian clusters intersect with heavy tails in this problematic region, which explains the pronounced confusion. A second troublesome area occurs around class 44 (“spider”). In the heat‑map visualization of stochastic block‑model embeddings (Fig. 5), a three‑class overlap can be observed, suggesting that the underlying feature distributions are not well separated. To alleviate these ambiguities, we construct a soft ensemble consisting of three 6464‑dimensional graph embeddings. The three vectors are concatenated and fed to a classifier trained with a standard softmax cross‑entropy loss. This strategy directly addresses the heavy‑tailed intersections observed in Fig. 7 by smoothing decision boundaries through complementary relational information. The empirical impact of the proposed approach on ImageNet‑100 is reported in Table 2, [5]. Compared with our hard‑embedding baseline, the soft ensemble raises Top‑1 accuracy from $82.70\%$ to $84.92\%$, while only modestly increasing model size (from $1.09\,\text{M}$ to $1.33\,\text{M}$ parameters) and computational cost (from $117.1\,\text{M}$ to $141.3\,\text{M}$ FLOPs). Although the soft‑embedding model is still considerably smaller than heavyweight architectures such as ResNet‑50 ($86.01\%$, 23.723.7M parameters) and Vision Transformer ($88.69\%$, 85.785.7M parameters), it attains a competitive accuracy–efficiency trade‑off.
Таблица 2: Top‑1 accuracy, number of parameters and FLOPs for various DNN models, [5]
Model ImageNet-100 Top‑1 Accuracy (%) Params FLOPs
Vgg16 82.07 134.7M 15.5G
MobileNetV2 82.60 2.4M 313.0M
MobileNetV3 83.15 4.3M 224.9M
CSPdarknet53 75.50 1.1M 159.5M
LightCSPNet dropout (p=0.2) 82.64 0.9M 143.9M
Hard Ensemble graph embedding (our) 82.70 1.09M 117.1M
Soft Ensemble graph embedding (our) 84.92 1.33M 141.3M
ResNet50 86.01 23.7M 4109.7M
Vision Transformer 88.69 85.7M 16.9G
@@ 5 Future Directions
The findings of this paper suggest several concrete research avenues. First, the present pipeline freezes the MobileNetV2 backbone, uses a cosine (or different kernel type) similarity kernel, and adopts a fixed MET‑QC‑LDPC adjacency; an obvious extension is to make all three components differentiable and optimize them jointly. One could learn the feature extractor $f_{\theta}$ together with the graph embedding, parameterize the similarity kernel (for example by learning a Mahalanobis metric) and back‑propagate through the adjacency construction, and finally adjust the circulant shift values that define the QC‑LDPC parity‑check matrix by gradient‑based or reinforcement‑learning methods. Rich multimodal embeddings such as those produced by Contrastive Language-Image Pre-training (CLIP) could serve as powerful initializations for this end‑to‑end training. A second line of inquiry is to embed topological information directly into the loss function. Since trapping sets correspond to unwanted $\mathbb{Z}/2$‑torsion and low‑genus cycles—manifested as negative eigenvalues of the Bethe–Hessian—it is natural to add differentiable penalties that discourage these structures. Possible regularizes include a term proportional to the number of negative Bethe–Hessian modes, an estimator of the first Betti number (for instance via a spectral trace approximation), or proxies for bordism obstructions such as the continuous genus $\widehat{A}$ or Stiefel–Whitney numbers. By penalizing these quantities during training we could suppress harmful subgraphs dynamically rather than removing them in a post‑hoc design step. Scaling the methodology to larger and more diverse data domains is another promising direction. Extending experiments from ImageNet‑10/100 to the full ImageNet‑10k benchmark will test whether the observed compression ratios and accuracy gains persist at million‑scale sample sizes. Moreover, applying the framework to video streams—where temporal coherence can be encoded as an additional edge type—and to multi-modal inputs (e.g., image–text pairs) by constructing heterogeneous MET‑QC‑LDPC graphs could reveal new ways of handling high‑dimensional, structured data while retaining a small memory footprint suitable for edge devices. The sparse, highly structured adjacency matrices produced by our graph design are also attractive as priors for modern Graph Neural Networks. Incorporating them into message‑passing GCN or GAT layers, using them to sparsify self‑attention mechanisms in Transformers, or employing diffusion‑based graph convolutions may combine the interpretability of physics‑inspired graphs with the expressive power of learned GNNs. An intriguing sub‑question is whether the Nishimori temperature $\beta_{N}$ can be treated as a learnable hyperparameter within such architectures. Finally, several theoretical questions remain open. It would be valuable to derive rigorous bounds linking reductions in Betti numbers (or other topological invariants) to improvements in classification margins, to explore connections between Ihara–Bass zeta function poles and persistent homology techniques from topological data analysis, and to generalize the quadratic‑Newton Nishimori estimator to other spin‑glass models such as Potts or vector spins that might better capture multi‑class structure. Addressing these issues will deepen our understanding of how statistical‑physics concepts translate into practical machine‑learning tools.